OpenStudy (anonymous):

1/2x^3 - x=0 explain please

OpenStudy (anonymous):

\[\frac{1}{2}x^3-x=0\] like that?

OpenStudy (anonymous):

\[x^3-2x=0\] \[x(x^2-2)=0\] \[x=0\] or \[x^2=0\] \[x=0,\sqrt{2},-\sqrt{2}\]